OpenStudy (anonymous):

[(30 + 6) − 32] ÷ 9 • 2 ANSWER Please I keep Getting It WRONG!!!!!

OpenStudy (anonymous):

You first do 30+6=36. Than you do 36-32=4. Next you divide by 9. \[4\div9=4/9\] Than you multiply by 2. 4/9 * 2=8/9. That's the answer I think.

OpenStudy (anonymous):

2/9? is that the answer? you have to remember the order of operations PEMDAS

OpenStudy (anonymous):

@sarah43 I think you are wrong. You have to do this from right to left. So you do like ([(30 + 6) − 32] ÷ 9) • 2 and not [(30 + 6) − 32] ÷ (9 • 2).

OpenStudy (anonymous):

yes i think ur rite srry